# BEAT: Balanced Frequency Adaptive Tuning for Long-Term Time-Series Forecasting

### v1

@@ 4.3 Ablation Study
5 Conclusion
This paper presents BEAT (Balanced frEquency Adaptive Tuning), a novel training framework that continuously monitors and adaptively adjusts the learning pace of each frequency sub-series. By dynamically modulating back-propagated gradients in accordance with real-time convergence feedback, BEAT ensures that both high-frequency and low-frequency model components converge in a balanced manner. Experimental results on a variety of benchmarks demonstrate that BEAT consistently outperforms state-of-the-art baselines, demonstrating the importance of resolving optimization conflicts due to mismatched learning speeds for different frequencies.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Detailed Architecture of Frequency-Specific Network
Figure 3: Illustration of Frequency-Specific Network
As we take WPMixer (Murad et al., 2025) as an instantiation for illustrating our proposed framework BEAT, the detailed structure of the Frequency-Specific Network is introduced as follows.
After the decomposition module breaks down the normalized time series into approximation and detail coefficient series, each wavelet coefficient series is transmitted through a separate resolution branch within the model, which consists of a RevIN normalization module, a patching and embedding module, multiple mixer modules, a head module, and a RevIN denormalization module. The total count of multivariate coefficient series or resolution branches in the model is determined as $f+1$ due to the presence of $f$ detail and $1$ approximation coefficient series.

### v2

@@ Introduction
Conclusion
In this paper, we explore a challenging problem in frequency-based time-series forecasting: the asynchronous convergence of high- and low-frequency sub-series. Although frequency-domain methods offer advantages for long-range forecasting, existing approaches overlook the mismatched learning dynamics across frequencies, leading to optimization conflicts and degraded performance. To tackle this, we propose BEAT (Balanced frEquency Adaptive Tuning), a framework that dynamically monitors the convergence of each frequency component and adaptively adjusts their gradient flow. By amplifying gradients for underfitting parts and suppressing overfitting ones, BEAT synchronizes learning across frequencies. Experiments on multiple benchmarks show that BEAT consistently improves long-range forecasting by resolving optimization conflicts caused by imbalanced learning speeds.